Dennis can eat a pie in 6 min. David can eat a pie in 12 min. Robert can eat a pie in 14.639145073927682623334797247841 min. If they share the same pie, how lomg will it take the three guys to scarf down the whole thing. (Hint: First figure out how many pies each guy can eat in an hour)
Answer by venugopalramana(3286)

Dennis can eat a pie in 6 min. David can eat a pie in 12 min. Robert can eat a pie in 14.639145073927682623334797247841 min. If they share the same pie, how lomg will it take the three guys to scarf down the whole thing. (Hint: First figure out how many pies each guy can eat in an hour)
DENIS CAN EAT IN 1 MINUTE....1/6 =0.166666667 PIE
DAVID CAN EAT IN 1 MINUTE....1/12 =0.083333333 PIE
ROBERT CAN EAT IN 1 MINUTE...1/14.63914507=0.06831 PIE
HENCE THE 3 TOGETHER CAN EAT IB 1 MINUTE=0.31831 PIE
HENCE THEY TAKE ...1/0.31831 = 3.14159153 MINUTES TO EAT THE PIE.